Three workers sort 66 boxes. Worker A sorts three times as many boxes as worker B, and worker B sorts two times as many boxes as worker C. How many boxes did worker B sort?
Answer by richwmiller(17219)

a=3b,
b=2c,
6c+2c+c=66
9c=66
3c=22
c=22/3= 7 1/3
b=44/3=14 2/3
a=44